OpenStudy (anonymous):

5.A bag contains 24 coloured counters, three blue, four red, nine green and eight yellow. If five counters are chosen at random from the bag, what is the probability of selecting exactly 3 green and no blue counters?

OpenStudy (kropot72):

The probability of selecting 3 green is given by \[P(3\ green)=\frac{9C3\times 15C2}{24C5}.......(1)\] The probability of selecting no blue is given by \[P(0\ blue)=\frac{21C5}{24C5}................(2)\] The required probability is found by multiplying (1) and (2).